Consider the parametric curve described by r(t) =〈2 sin(3t), √4t, 2cos(3t)〉. Find the unit tangent vector of r(t).
1
Expert's answer
2020-05-25T15:56:40-0400

We have given the parametric curve (vector valued) function as

"r(t)=\big<2\sin(3t),\sqrt{4t},2\cos(3t)\big>"

Since, we know that tangent vector to "r(t)" is given

"\frac{dr(t)}{dt}=r'(t)"

Thus,

"r'(t)=\big<\frac{d(2\sin(3t))}{dt},\frac{d(\sqrt{4t})}{dt},\frac{d(2\cos(3t))}{dt}\big>\\
\implies r'(t)=\big<6\cos(3t),\frac{1}{\sqrt{t}},-6\sin(3t)\big>"

We also know that, unit vector for any vector "v" is

"\hat{v}=\frac{v}{||v||}"

where, "||v||" is the norm of "v" . Hence norm of tangent vector of "r(t)" will be

"||r'(t)||=\sqrt{(6\sin(3t))^2+(\frac{1}{\sqrt{t}})^2+(-6\cos(3t))^2}
\\
\implies ||r'(t)||=\sqrt{36+\frac{1}{t}}"

Therefore,

"\widehat{r'(t)}=\frac{r'(t)}{||r'(t)||}\\
\implies \widehat{r'(t)}=\big<\frac{6\cos(3t)}{\sqrt{36+\frac{1}{t}}},\frac{1}{\sqrt{t}\sqrt{36+\frac{1}{t}}},\frac{-6\sin(3t)}{\sqrt{36+\frac{1}{t}}}\big>\\
\implies \widehat{r'(t)}=\big<\frac{6\cos(3t)}{\sqrt{36+\frac{1}{t}}},\frac{1}{\sqrt{36t+1}},\frac{-6\sin(3t)}{\sqrt{36+\frac{1}{t}}}\big>"

Hence the required unit vector is

"\big<\frac{6\cos(3t)}{\sqrt{36+\frac{1}{t}}},\frac{1}{\sqrt{36t+1}},\frac{-6\sin(3t)}{\sqrt{36+\frac{1}{t}}}\big>"

We are done.